2) An automobile manufactured claims that a product will, starting from rest, travel 0.40 km in nine seconds...
3) A certain car is capable of accelerating at a rate of .50 m/s squared, How long does it take for this car to go from the speed of 55 KM/h to speed of 60 KM/h

Please help I’ve tried multiple times and don’t know what I’m doing wrong!
(2) An automobile manuPacturer claims that is product usil,s kem rest, trave 4o km in 9.os what is the magnitu constont acceleratton reqvtred to do-this ? and (3/) A certoin car is capable of acceleratin, at a rate ap .50 msH long does it take fo tas aor to g0 Speed of 53 Km/h to a spech of GoKm/h?
Answer #1

1) Given Distance S = 0.4km = 400 m

Time t = 9 s

Then:

S =ut + rac{1}{2}at^2, Since the car starts from rest, u = 0 m/s

400 =-a(9)

× 400-0876m 92

2) Initial Speed = 55 km/h = 15.27 m/s

Final Speed = 60 km/h = 16.67 m/s

Then v = u + at

16.67 = 15.27 + 0.5 t

t = 2.8 s

( Note: To convert km/h to m/s , multiply by 5/18 )